Use assembly language to

write a complete program that will input values fornum1 ,num2 and num3 and display the

value of the expression ( (num1 ^ 3) * num2 + 5 * ( num2 ^ 2) ) / num3.

assume that the user enters only numbers that are greater than zero and the calculation never exceed 4 bytes size.

Sample run:

num1 = 1
num2 = 2
num3 = 3
((num1 ^ 3) * num2 + 5 * ( num2 ^ 2)) / num3 = 7 R 1
Answer #1

Here is the assembly code to calculate the value of the expression ( (num1 ^ 3) * num2 + 5 * ( num2 ^ 2) ) / num3.

Where the numbers are greater than zero & size of calculation does not exceeds 4 bytes.

org 0000h ;sets the program counter to the 0000h

mov R0,#num1 ;assign value 'num1' to R0 which is converted to its equivalent hexadecimal value

mov R1,#num2 ;assign value 'num2' to R1 which is converted to its equivalent hexadecimal value

mov R2,#num3 ;assign value 'num3' to R2 which is converted to its equivalent hexadecimal value

mov A,R0 ;mov the value of R0 to the accumulator A means A=num1

mul A R0 ;multiple the values of A and Ro and stores result in A,A=num1*num1

mul A R0 ;again multiple the values of A and Ro and stores result in A,A=num1*num1*num1

mul A R1 ;multiply value of A with value of R1,A=(num1^3)*num2

mov R3,A ;store the result of A in R3

mov A,R1 ;put the value of num2 to A

mul A R1 ;multiply values of A and R1,A=num2*num2

mul A #5 ;multiply the value of A with 5,A=(num2^2)*5

mov R4,A ;store the result of A in R4

mov A,R3 ;put the value of R3 to A

add A,R4 ;add the value of R4 with A & store addition in A

div A,R2 ;divide the value of A with R2 i.e. num3

end